solve using any method
2x^2=8x

Answer by checkley77(12844)

2x^2=8x
2X^2-8X=0
X(2X-8)=0
X=0
2X-8=0
2X=8
X=8/2
X=4 ANS.
PROOF:
2(4)^2=8*4
2*16=32
32=32